# Hacking Task Confounder in Meta-Learning

### v1

@@ Introduction
To explore the reasons behind the task confounders mentioned above, we construct an SCM for meta-learning regarding two tasks $\tau_{i}$ and $\tau_{j}$ based on data generation mechanisms, as illustrated in Figure 1(a). In this model, $Y_{i}$ and $Y_{j}$ denote the label variables for tasks $\tau_{i}$ and $\tau_{j}$ respectively, while $X_{i}$ and $X_{j}$ signify the sample variables for the two tasks, respectively. Additionally, ${{\rm A}^{i}}$ and ${{\rm A}^{j}}$ represent distinct sets of causal factors exclusive to tasks $\tau_{i}$ and $\tau_{j}$, respectively, such as color, shape, and texture. On the other hand, ${{\rm B}^{i,j}}$ encompasses causal factors shared between tasks $\tau_{i}$ and $\tau_{j}$. We assume that the samples and the labels are both generated by a ground-truth causal mechanism following (Suter et al. 2019; Hu et al. 2022). Specifically, we assume that the sample is generated by disentangled causal mechanisms, e.g., $p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}})=\prod\nolimits_{k}{p({X_{i}}|{{\rm A}^{i}_{k}})}\prod\nolimits_{t}{p({X_{i}}|{\rm B}_{t}^{i,j})}$, where ${\rm A}^{i}_{k}$ denotes the $k$-th element of ${\rm A}^{i}$ and ${\rm B}_{t}^{i,j}$ denotes the $t$-th element of ${\rm B}^{i,j}$. As ${\rm A}^{i}$, ${\rm A}^{j}$, and ${\rm B}^{i,j}$ represent high-level knowledge of the data, we could naturally define the task label variable $Y_{i}$ for task $i$ as the cause of the ${\rm B}^{i,j}$ and ${\rm A}^{i}$. For the task $\tau_{i}$, we call ${\rm B}^{i,j}$ and ${\rm A}^{i}$ as the causal feature variables that are causally related to $Y_{i}$, and we call ${\rm A}^{j}$ as the non-causal feature variables to task $\tau_{i}$. Therefore, we have $p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}},{{\rm A}^{j}})=p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}})$.
@@
Based on the proposed SCM, an ideal meta-learning predictor for each task should only utilize causal factors and be invariant to any intervention on non-causal factors. However, the joint learning of multiple tasks can give rise to the issue of spurious correlations, thereby making it challenging to achieve optimal predictions. In order to investigate the mechanisms underlying the generation of spurious correlations, we consider the scenario of two binary classification tasks. Let $Y_{i}$ and $Y_{j}$ be variables from {±$\left\{{\pm 1}\right\}$. We assume the two tasks have non-overlapping factors, e.g., ${{\rm{B}}^{i,j}}=\emptyset$, and the elements in ${{\rm A}^{i}}$ and ${{\rm A}^{j}}$ satisfy the constraint of Gaussian distribution. Then, we have:
Theorem 1
@@
If the correlation between $Y_{i}$ and $Y_{j}$ is not equal to 0.5, the optimal classifier has non-zero weights for non-causal factors for each task. If the correlation between $Y_{i}$ and $Y_{j}$ equals 0.5 and the number of training samples is limited, the optimal classifier also has non-zero weights for non-causal factors for each task.
@@
As inferred from the aforementioned theorem, the learned model leverages the causal factors from other tasks to facilitate the learning of the target task. The SCM corresponding to this process is illustrated in Figure 1(b). Taking the task $\tau_{i}$ as an example, the meta-learning model uses the causal factors ${{\rm A}^{j}}$ belonging to the task $\tau_{j}$ for learning $Y_{i}$. Thus, there is a spurious correlation between ${{\rm A}^{j}}$ and $Y_{i}$, which can be represented as a spurious path ${{\rm A}^{j}}\to Y_{i}$. Also, we can obtain the spurious path ${{\rm A}^{i}}\to Y_{j}$. The learning process can be viewed as the inverse process of the generating mechanism. Therefore, we can obtain the SCM with two spurious paths, which can reflect the internal mechanism of task confounders in multi-task learning. The proof of Theorem 1 is provided in Appendix A.
Methodology
@@
Based on the above evidence and analysis, we can know that task confounders can cause spurious correlations between causal factors and labels. An ideal meta-learning model should learn multi-task knowledge in a shared representation and identify which portions of knowledge are causally related to each task. This inspires us to propose a method called MetaCRL that can encode decoupled causal knowledge. It serves as a plug-and-play learner that consists of two modules: the disentangling module and the causal module. The disentangling module aims to acquire all causal factors and provide subsets of causal factors specifically relevant to individual tasks, while the causal module aims to ensure the causality of factors in the disentangling module. The pseudocode of our MetaCRL is provided in Appendix B.
Disentangling Module
For a pre-trained CNN-based encoder, each channel can be regarded as being related to a kind of semantic information or visual concept (Islam, Jia, and Bruce 2020). Thus, we can use the data representation to learn the causal factors. During the training phase, we denote the $N_{tr}$ training tasks as $\left\{\tau_{i}\right\}_{i=1}^{N_{tr}}$. Suppose that the number of causal factors is $N_{k}$, then, we propose obtaining these $N_{k}$ factors through the learning of a matrix $\Xi\in{{\mathbb{R}}^{N_{z}\times N_{k}}}$, where $N_{z}$ represents the dimension of the feature representation, e.g., the output dimension of the encoder $g$, and each column of $\Xi$ represents a distinct factor. Based on $\Xi$, we can obtain a new representation of each sample, which can be called a causal representation, e.g., the causal representation for $x_{i,j}^{s}$ can be presented as ${\Xi^{\rm{T}}}g(x_{i,j}^{s})$. Based on the causal analysis above, the causal factors should be divided into $N_{tr}$ overlapping groups, and each group corresponds to a task. To obtain these groups, we propose a learnable grouping function $f_{gr}$, which is implemented using Multi-Layer Perceptrons (MLPs). Specifically, for task $\tau_{i}$, we first calculate the average sample $x_{i}$ for this task, e.g., ${x_{i}}=\frac{1}{N_{i}^{s}+N_{i}^{q}}(\sum\nolimits_{j=1}^{N_{i}^{s}}{x_{i,j}^{s}}+\sum\nolimits_{j=1}^{N_{i}^{q}}{x_{i,j}^{q}})$. Subsequently, $x_{i}$ is input into $g$, $\Xi$, and $f_{gr}$, e.g., ${f_{gr}}({\Xi^{\rm{T}}}g({x_{i}}))$, yielding a vector with all elements greater than zero and matching the dimensionality of the causal representation. Concurrently, each element is subject to the normalization operation $\rm{Norm}$. As a result, the individual elements of the output vector ${\rm{Norm}}(f_{gr})$ can be interpreted as the probabilities that each causal factor belongs to $\tau_{i}$. In a word, we can obtain the causal factors for each task based on $\tau_{i}$ and $f_{gr}$.
Ideally, each factor in $\Xi$ should represent a distinct semantic, and changing one factor should not affect others. However, in reality, without explicit constraints, the factors in $\Xi$ may still be correlated, hindering the learning of causal structures (Locatello et al. 2019). Therefore, we propose a regularization term to achieve complete decoupling. Specifically, we directly penalize the similarity between different factors, which can be formulated as:
@@
where ${{\Xi_{:,i}}}$ represents the $i$-th column of $\Xi$. As we can see, minimizing $\mathcal{L}_{\rm{DM}}(\Xi)$ can lead to a similarity of 0 between different causal factors, thereby empirically establishing independence among distinct factors.
Note that each task is only related to a small number of factors, and the factors of different tasks can vary greatly. It may lead to degenerate solutions in which only a few factors are utilized. Therefore, we need to constrain the output of $f_{gr}$ to be sparse and diverse. To achieve this, we propose to use the $L_{1}$ norm with an entropy term to the output of $f_{gr}$:
@@
where ${{f_{gr}}{{({\Xi^{\rm{T}}}g({x_{i}}))}_{j}}}$ represents the $j$-th element of the output of $f_{gr}$. We can see that minimizing the first term of $\mathcal{L}_{\rm{DM}}(f_{gr})$ can make the output of $f_{gr}$ to be sparse, and maximizing the second term of $\mathcal{L}_{\rm{DM}}(f_{gr})$ can make the output of $f_{gr}$ to be diverse.
By combining Eq.3 and Eq.4, the loss of the disentangling module can be presented as:
@@
where $\lambda_{1}$ and $\lambda_{2}$ denote the loss weights of $\mathcal{L}_{\rm{DM}}(\Xi)$ and $\mathcal{L}_{\rm{DM}}(f_{gr})$, respectively.
Causal Module
In the disentangling module, each column of $\Xi$ is treated as a causal factor. However, due to the random initialization of matrix $\Xi$ and the absence of explicit constraints in the disentangling module, ensuring causality in $\Xi$ becomes challenging. Moreover, the causal factors for each task also exhibit randomness, making it difficult to guarantee their authenticity. To solve these problems, we propose the causal module to ensure causality in the disentangling module. Following (Koyama and Yamaguchi 2020), we can see that a model invariant to multiple distributions could learn causal correlations. Also, based on Theorem 9 described in (Arjovsky et al. 2019), by enforcing invariance over multiple training datasets that exhibit distribution shifts, the task-specific models should only utilize causal factors that are helpful to themselves, and assign zero weights to those non-causal factors for the task. Therefore, the causal module is designed to facilitate causal learning by using this invariance.
As widely known, during the training phase of meta-learning, the training data can be divided into multiple support sets and multiple query sets. As they comprise different samples, they can be regarded as distinct data distributions with distributional shifts. Meanwhile, the learning process in Eq.2 can be depicted as follows: First, for every $f_{\theta}$, optimizing Eq. 1 can achieve an optimal $f_{\theta}^{i}$ and $\mathcal{L}(Y_{i}^{s},X_{i}^{s},f_{\theta}^{i})$ on the support set. Subsequently, altering the value of $f_{\theta}$ impacts the optimal $f_{\theta}^{i}$ accordingly. At last, seek the optimal $f_{\theta}$ to obtain the optimal $f_{\theta}^{i}$ that can optimize $\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathcal{L}(Y_{i}^{q},X_{i}^{q},f_{\theta}^{i})$ on the query set. Consequently, the optimization of Eq.1 and Eq.2 can be interpreted as achieving optimality across multiple datasets using the same $f_{\theta}$. Based on the above illustration, we propose to utilize a bi-level optimization mechanism to learn $\Xi$ and $f_{gr}$, e.g., similar to Eq.1 and Eq.2, thus ensuring causality. Specifically, for the first level, we learn $\Xi^{{}^{\prime}}$ and ${f}_{gr}^{{}^{\prime}}$ with the support sets through the following objectives:
@@
where $\odot$ represents the element-wise multiplication operator between two vectors, while $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ are the learning rates. It’s worth noting that $\mathcal{L}(Y_{i}^{\cdot},X_{i}^{\cdot},\Xi,{f_{gr}})$ is optimized using the causal representation ${\Xi^{\rm{T}}}g(x_{i,j}^{\cdot})$ with the weight ${\rm{Norm[}}{f_{gr}}({\Xi^{\rm{T}}}g({x_{i}}))]$. The weight is intended to restrict the causal features of samples in the task $\tau_{i}$ to be associated only with a subset of causal factors.
The learning process of $\Xi$ and $f_{gr}$ can be regarded as enforcing invariance over the support sets and the query sets, thus, the bi-level optimization mechanism for $\Xi$ and $f_{gr}$ can ensure causality. Meanwhile, Eq.6 and Eq.7 are learned based on pre-defined $h$ and $g$, thus rendering the MetaCRL a plug-and-play learner.
Overall Objective
In this subsection, we embed the above causal representation learning process into a meta-learning framework for joint optimization. The training process for MetaCRL in each batch is divided into two steps. In the first step, with $\Xi$ and $f_{gr}$ held fixed, we optimize $h$ and $g$. Specifically, the objective of the inner loop mentioned in Eq.1 becomes:

### v4

@@ 3.3 Causal Analysis and Motivation
To explore the reasons behind the above phenomenon, we propose using causal theory for analysis. We first construct a Structural Causal Model (SCM) based on the ground-truth causal mechanisms  Suter et al. (2019); Hu et al. (2022), as shown in Figure 2(a). Specifically, this SCM contains two tasks $\tau_{i}$ and $\tau_{j}$, where $Y_{i}$ and $Y_{j}$ denote the label variables for tasks $\tau_{i}$ and $\tau_{j}$, $X_{i}$ and $X_{j}$ signify the corresponding generated samples for these two tasks, respectively. Meanwhile, ${{\rm A}^{i}}$ and ${{\rm A}^{j}}$ represent the distinct sets of causal factors specific to tasks $\tau_{i}$ and $\tau_{j}$, while ${{\rm B}^{i,j}}$ encompasses shared causal factors. In this SCM, we assume that the samples $X_{i}$ and $X_{j}$ are generated by disentangled causal mechanisms using the causal factors, then $p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}})=\prod\nolimits_{k}{p({X_{i}}|{{\rm A}^{i}_{k}})}\prod\nolimits_{t}{p({X_{i}}|{\rm B}_{t}^{i,j})}$, where ${\rm A}^{i}_{k}$ denotes the $k$-th factor of ${\rm A}^{i}$, and ${\rm B}_{t}^{i,j}$ denotes the $t$-th factor of ${\rm B}^{i,j}$. Since ${\rm A}^{i}$, ${\rm A}^{j}$, and ${\rm B}^{i,j}$ represent high-level knowledge of the data, we could naturally define the task label variable $Y_{i}$ for task $i$ as the cause of the ${\rm B}^{i,j}$ and ${\rm A}^{i}$. For the task $\tau_{i}$, we call ${\rm B}^{i,j}$ and ${\rm A}^{i}$ as the causal feature variables that are causally related to $Y_{i}$, and we call ${\rm A}^{j}$ as the non-causal feature variables to task $\tau_{i}$. Therefore, we have $p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}},{{\rm A}^{j}})=p({X_{i}}|{{\rm A}^{i}},{{\rm{B}}^{i,j}})$.
@@
Based on the proposed SCM, an ideal meta-learning predictor for each task should only utilize causal factors and be invariant to any intervention on non-causal factors. However, the joint learning of multiple tasks in meta-learning could give rise to the issue of using non-causal factors for unseen tasks, also known as spurious correlations, thereby making it challenging to achieve optimal predictions. To verify this claim, we consider the scenario of two binary classification tasks for simple but clear explanations. Let $Y_{i}$ and $Y_{j}$ be variables from {±$\left\{{\pm 1}\right\}$, we assume $\tau_{i}$ and $\tau_{j}$ have non-overlapping factors, i.e., ${{\rm{B}}^{i,j}}=\emptyset$, and the elements in ${{\rm A}^{i}}$ and ${{\rm A}^{j}}$ satisfy the constraint of Gaussian distribution. Then, we have:
Theorem 1.
@@
If the correlation between $Y_{i}$ and $Y_{j}$ is not equal to 0.5, the optimal classifier has non-zero weights for non-causal factors for each task. If the correlation between $Y_{i}$ and $Y_{j}$ equals 0.5 with limited training data, the optimal classifier also has non-zero weights for non-causal factors in each task.
@@
As inferred from the aforementioned theorem, the learned model leverages the causal factors from other tasks to facilitate the learning of the target task. Taking the task $\tau_{i}$ as an example, the meta-learning model uses the causal factors ${{\rm A}^{j}}$ belonging to the task $\tau_{j}$ for learning $Y_{i}$. Therefore, there is a spurious correlation between ${{\rm A}^{j}}$ and $Y_{i}$, which can be represented as a spurious path ${{\rm A}^{j}}\to Y_{i}$. Similarly, we can obtain the spurious path ${{\rm A}^{i}}\to Y_{j}$ for task $\tau_{j}$. These spurious correlations are called “task confounders”, which are the reasons that lead to negative knowledge transfer in Subsection 3.2. The learning process can be viewed as the inverse process of the generating mechanism. Therefore, we can obtain the SCM with two spurious paths as illustrated in Figure 2(b), which reflects the internal mechanism of task confounders in multi-task learning. The proof is provided in Appendix A.
4 Methodology
@@
Based on the above analysis, we know that task confounders cause spurious correlations between causal factors and labels. An ideal meta-learning model should identify knowledge that is causally related to each task and learn from the identified multi-task knowledge. Therefore, we propose MetaCRL, a plug-and-play meta-learning causal representation learner that can encode decoupled causal factors for more efficient ML. It consists of two modules: (i) the disentangling module which aims to extract generating factors and eliminate task confounders; and (ii) the causal module which aims to ensure the causality of the obtained generating factors. In this section, we first introduce the disentangling module and the causal module in Subsections 4.1 and 4.2, respectively. Next, we provide the overall objective in Subsection 4.3. The pseudocode and pipeline of MetaCRL are shown in Appendix B.
4.1 Disentangling Module
In this module, we aim to obtain the whole generating factors related to all tasks and the task-specific generating factors related to each single task. Specifically, we first obtain the whole generating factors by learning a semantic matrix $\Xi$. Next, we use a grouping function $f_{gr}$ to acquire subsets of generating factors relevant to every single task. Note that this module does not guarantee the causality of the obtained generating factors, which will be addressed in the causal module.
For a pre-trained encoder, different channels of the feature representations are related to different kinds of semantics Islam et al. (2020). Thus, we propose to use the feature representation to learn the generating factors. During the training phase, we denote the $N_{tr}$ training tasks as $\left\{\tau_{i}\right\}_{i=1}^{N_{tr}}$. Suppose that the number of generating factors is $N_{k}$, then, we propose obtaining these $N_{k}$ factors through the learning of a matrix $\Xi\in{{\mathbb{R}}^{N_{z}\times N_{k}}}$. Here, $N_{z}$ represents the dimension of the feature representation, i.e., the output dimension of the encoder $g$, and each column of $\Xi$ represents a distinct factor. Based on $\Xi$, we can obtain a new representation of each sample, which can be called a generating representation, e.g., the generating representation for $x_{i,j}^{s}$ can be presented as ${\Xi^{\rm{T}}}g(x_{i,j}^{s})$.
Generally, generating factors in geometric space can be conceptualized as coordinate basis vectors, where each generating factor corresponds to a specific basis vector Jensen and Shen (2004). Moreover, different coordinate bases can undergo mutual transformations via a reversible matrix, implying their equivalence. Hence, learning a task-specific matrix, serving as a base matrix, allows us to approximate task-related generating factors. Therefore, for $\Xi$ to be considered a generating factor matrix, we need to constrain the column vectors of $\Xi$ to be orthogonal to each other. Then we have:
@@ 4.1 Disentangling Module
where ${{\Xi_{:,i}}}$ represents the $i$-th column of $\Xi$. Minimizing $\mathcal{L}_{\rm{DM}}(\Xi)$ makes the different columns of $\Xi$ orthogonal to each other, thus leading $\Xi$ to be task-related generating factors.
Next, for all the $N_{tr}$ training tasks, the generating factors should be divided into $N_{tr}$ overlapping groups, and each group corresponds to a task. To obtain these groups, we propose a learnable grouping function $f_{gr}$, which is implemented using Multi-Layer Perceptrons (MLPs) to acquire task-specific generating factors. Take task $\tau_{i}$ as an example, we first calculate the average sample $x_{i}$ for this task, i.e., ${x_{i}}=\frac{1}{N_{i}^{s}+N_{i}^{q}}(\sum\nolimits_{j=1}^{N_{i}^{s}}{x_{i,j}^{s}}+\sum\nolimits_{j=1}^{N_{i}^{q}}{x_{i,j}^{q}})$. Then, we input $x_{i}$ into the encoder $g$, $\Xi$, and $f_{gr}$, i.e., ${f_{gr}}({\Xi^{\rm{T}}}g({x_{i}}))$, yielding a vector with all elements greater than zero and matching the dimensionality of the generating representation. Then, each element is subject to the normalization operation, denoted as $\rm{Norm}(\cdot)$. As a result, the individual elements of the output vector, i.e., ${\rm{Norm}}(f_{gr})$, can be interpreted as the probabilities that each generating factor belongs to task $\tau_{i}$.
Note that each task is associated with a subset of factors in $\Xi$ and can vary significantly from task to task. Meanwhile, the above calculation process of $\Xi$ and $f_{gr}$ may lead to degenerate solutions, e.g., the subset of generating factors for each task is the same. To address this issue, we propose a regularization term that consists of a $L_{1}$ norm and an entropy term, constraining the output of $f_{gr}$ to be sparse and diverse. By minimizing the $L_{1}$ norm, we make the output of $f_{gr}$ sparse, ensuring obtain subsets of generating factors only relevant to each single task. By maximizing the entropy term, we make the output of $f_{gr}$ diverse, preventing the acquisition of task-specific generating factors suffering degenerate solutions. The regularization term is:
@@
where ${{f_{gr}}{{({\Xi^{\rm{T}}}g({x_{i}}))}_{j}}}$ represents the $j$-th element of the output of $f_{gr}$. Through Eq.5, we obtain accurate task-specific generating factors, thus eliminating task confounders.
By combining Eq.4 and Eq.5, we obtain the loss of the disentangling module which can be expressed as:
@@
where $\lambda_{1}$ and $\lambda_{2}$ denote the loss weights of $\mathcal{L}_{\rm{DM}}(\Xi)$ and $\mathcal{L}_{\rm{DM}}(f_{gr})$, respectively. Through the above process with three constraints, i.e., correlation, sparsity, and diversity, we can accurately obtain all the generating factors and the task-specific generating factors without task confounders.
4.2 Causal Module
In this module, we aim to ensure the causality of the generating factors obtained in the disentangling module. Following  Koyama and Yamaguchi (2020), a model invariant to different distributions can learn causal correlations. Meanwhile, based on Theorem 9 described in Arjovsky et al. (2019), by enforcing invariance over multiple training datasets that exhibit distribution shifts, the task-specific models could only use task-related causal factors and assign zero weights to those non-causal generating factors. Therefore, the causal module is designed to facilitate causal learning by using this invariance, thereby ensuring the causality of the generating factors obtained by $\Xi$ and $f_{gr}$.
During the training phase of ML, the training data can be divided into multiple support sets and query sets. As they comprise different samples, they can be regarded as different data distributions with distributional shifts. Meanwhile, the learning process of meta-learning can be depicted as follows: First, for every $f_{\theta}$, optimizing Eq.1 can achieve an optimal $f_{\theta}^{i}$ and $\mathcal{L}(Y_{i}^{s},X_{i}^{s},f_{\theta}^{i})$ on the support set. Next, altering the value of $f_{\theta}$ impacts the optimal $f_{\theta}^{i}$, we seek the optimal $f_{\theta}$ to obtain the optimal $f_{\theta}^{i}$ by optimizing $\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathcal{L}(Y_{i}^{q},X_{i}^{q},f_{\theta}^{i})$ on the query sets (Eq.2). Thus, the bi-level optimization of Eq.1 and Eq.2 can be interpreted as achieving optimality across multiple datasets using the same $f_{\theta}$, and the causal factors are invariant on the support and query sets of the same task.
Based on the above illustration, we propose to utilize a bi-level optimization mechanism to learn $\Xi$ and $f_{gr}$ which is similar to Eq.1 and Eq.2, thus ensuring causality. Specifically, for the first level, we learn $\Xi^{{}^{\prime}}$ and ${f}_{gr}^{{}^{\prime}}$ with the support sets through the following objectives:
@@ 4.2 Causal Module
where $\odot$ represents the element-wise multiplication operator between two vectors, i.e., the generating representation ${\Xi^{\rm{T}}}g(x_{i,j}^{\cdot})$ and the weight ${\rm{Norm[}}{f_{gr}}({\Xi^{\rm{T}}}g({x_{i}}))]$, while $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ and $\alpha_{4}$ are the learning rates. Note that both in Eq.7 and Eq.8, the loss $\mathcal{L}(Y_{i}^{\cdot},X_{i}^{\cdot},\Xi,{f_{gr}})$ is calculated using the generating representations with causal weights instead of feature representations, which restrict the features of the samples in $\tau_{i}$ to be associated only with task-specific causal factors.
In summary, the learning process of $\Xi$ and $f_{gr}$ can be regarded as enforcing invariance over the support sets and the query sets, and the bi-level optimization mechanism for $\Xi$ and $f_{gr}$ can ensure causality. Meanwhile, $\Xi$ and $f_{gr}$ are learned independently with the fixed meta-learning model $f_{\theta}$ in the middle training following modularity design, thus rendering the MetaCRL a plug-and-play learner.
4.3 Overall Objective
In this subsection, we embed the above causal representation learning process into a meta-learning framework for joint optimization. The training process with MetaCRL in each batch is divided into two steps. In the first step, with $\Xi$ and $f_{gr}$ held fixed, we optimize the meta-learning model $f_{\theta}=h\circ g$. Specifically, the objective of the inner loop becomes: